OpenStudy (anonymous):

the equation of a line is y=-1/2x-1. what is the equation of the line that is perpendicular to the first line and passes through the point (2,-5)?

OpenStudy (anonymous):

y = -x/2 - 1 So, it's slope is -1/2. The slope of the line perpendicular to this one will be 2. Because the product of the slopes of perpendicular lines must be -1. Now, you can write the new line as: y = 2x + b. Because it passes through x=2 and y=-5, you can substitute those values to find b. Then you will have your line equation.

OpenStudy (anonymous):

i dont understand these are my ansers: A.y=2x+1 B.y=1/2X-6 C.y=-2x-1 d.y=2x-9

OpenStudy (mathmate):

The same question has been answered in your other post.